Find the domain of the following function. X +3 f(x)= x2 - 49
Answer #1

fcx) - I + 3 12.49 z?- 49 = → x² - 72 = = (x++) (1-7) 1N OC = -7) 3 - 7 Domain: AU R except x = 7,-7